# Bohao Wu   Qingyun Wang   Yue Guo

### v2

@@ Abstract
Personalizing jargon detection and explanation is essential for making technical documents accessible to readers with diverse disciplinary backgrounds. However, tailoring models to individual users typically requires substantial annotation efforts and computational resources due to user-specific finetuning. To address this, we present a systematic study of personalized jargon detection, focusing on methods that are both efficient and scalable for real-world deployment. We explore two personalization strategies: (1) lightweight fine-tuning using Low-Rank Adaptation (LoRA) on open-source models, and (2) personalized prompting, which tailors model behavior at inference time without retaining. To reflect realistic constraints, we also investigate hybrid approaches that combine limited annotated data with unsupervised user background signals. Our personalized LoRA model outperforms GPT-4 by 21.4% in F1 score and exceeds the best performing oracle baseline by 8.3%. Remarkably, our method achieves comparable performance using only 10% of the annotated training data, demonstrating its practicality for resource-constrained settings. Our study offers the first work to systematically explore efficient, low-resource personalization of jargon detection using open-source language models, offering a practical path toward scalable, user-adaptive NLP system.  [^1] Code can be found at: anonymous GitHub repository.
@@ 1 Introduction
Figure 1: Bridging disciplinary gap with limited compute and annotation: Our approach leverages PEFT, background-aware prompting, and mixed supervision to enable efficient and effective model adaption.
@@
Figure 2: To enable personalized jargon detection, we first fine tune a LoRA based PEFT model using supervised, unsupervised, and hybrid training strategies that reflect real world scenarios with varying levels of annotation availability. Next, we enhance the contextual understanding of the target researcher through a range of background aware prompting methods, including vanilla, metadata based, profile enhanced, and two oracle prompting configurations using nearest neighbor techniques (researcher and abstract), to generate personalized familiarity predictions.
@@
Our work builds on Guo et al. (2024), who introduced the first benchmark with scholar annotations for interdisciplinary jargon and explored the use of GPT-4 OpenAI (2023) for personalized jargon detection. While innovative, their method relies on computationally expensive LLM prompting and requires access to detailed metadata and annotated examples, limiting scalability, generalization, and applicability in low-resource or cold-start settings. We propose a more efficient and generalizable solution using open-source LLMs and unsupervised PEFT techniques. Our contributions are:
  • Lightweight Personalized PEFT: Introducing the first LoRA-based PEFT method for personalized jargon detection that outperforms prior work while significantly reducing computational overhead.
  • Background-Aware Prompting: Designing personalized prompting strategies using researcher background, yielding efficient performance gains across domains.
  • Unsupervised Personalization: Integrating unsupervised user data achieves strong performance with only 10% labeled supervision, enhancing accessibility and generalizability.
@@ 2 Method
To model each researcher’s background, we first retrieve contextual information §2.2 and explore three fine-tuning settings: supervised, unsupervised, and hybrid §2.3.
@@ 2.1 Experimental Setup
We use the personalized jargon detection dataset from Guo et al. (2024), which contains 11k term familiarity (i.e., familiar or unfamiliar) annotations provided by 11 computer science researchers for terms extracted from 100 paper abstracts. To our knowledge, it is the only interdisciplinary jargon dataset with high-quality annotations, rich personal metadata, and accompanying published papers from the annotators. We follow the same data split as Guo et al. (2024), dividing the dataset into 60/20/20 for the train, validation, and test sets. We selected Llama-3.1 8B Instruct 4bit Dubey et al. (2024) as our baseline model for personalized jargon detection based on its strong instruction-following ability, low mismatch rate, and competitive performance in preliminary evaluations across several state-of-the-art LLMs. We evaluate model performance using the effective F-1 Score, which adjusts for output format errors by penalizing predictions with a high mismatch rate. Details in App. §A.
@@ 2.2 Background-Aware Prompting
Building on prior work Guo et al. (2024); Tan et al. (2024b), we design a set of prompting strategies that incorporate varying levels of researcher-specific context into the model input. These strategies differ in both the type and richness of the background information provided:
  • Metadata: Structured features including the annotator’s self-defined subfield (e.g., NLP, computer vision), publication count, average references, year of first publication, and the domain of the current abstract. These serve as lightweight indicators of expertise and familiarity.
  • Profile-enhancement: Following prior work on user modeling in personalized recommendation Tan et al. (2024b), we use the baseline model to generate a natural language summary of the annotator’s research background based on their metadata.
To explore upper-bound performance, we also include two oracle prompting configurations based on social recommendation Kang et al. (2022); Guy et al. (2009); Goldberg et al. (1992):
@@
As a baseline, we include a vanilla prompting setup in which no contextual information about the annotator is provided.
2.3 Lightweight Personalized PEFT
  • Supervised: We adopt LoRA (Hu et al., 2022) for PEFT, following findings from Tan et al. (2024b) demonstrating its strong performance. In this setting, the Instruction field includes background context retrieved as described in §$2.$
  • Unsupervised: To simulate low-resource scenarios, we fine-tune using only unlabeled data, including titles and abstracts from each annotator’s previous publications. For annotators with limited publication history, we supplement the data with papers from their self-defined subdomain. This reflects practical cases where personalized models must be trained on domain-relevant but unlabeled content.
  • Hybrid: To examine the trade-off between annotation effort and personalization quality, we also explore a hybrid approach that combines a small amount of labeled data with the unlabeled corpus. This setting aims to inform future annotation strategies by evaluating how much supervision is sufficient for reasonable model performance.
@@ 3 Results
Figure 3: Validation performance of supervised (a), unsupervised (b), and hybrid (c) models in jargon familiarity detection. Each point represents a model fine tuned for one of the 11 annotators using various prompting strategies. GPT-4 and Oracle baselines indicate prior state-of-the-art performance from Guo et al. (2024). For hybrid models, training steps are used on the x-axis to enable consistent comparison across datasets of different sizes and to reflect relative computational cost.
Supervised fine-tuning outperforms GPT-4 with strong gains across prompting strategies
Fig. 3 (a) shows the validation performance of our supervised model trained on the full annotated dataset across different prompting strategies and training epochs. All methods show rapid performance gains in the first 20 epochs, after which improvements plateau. Among prompting strategies, vanilla prompting achieves the highest F1 score of 0.82, suggesting that even without explicit background context, the model benefits significantly from full supervision. Metadata-enhanced and profile-enhanced prompts yield comparable results, both saturating around 0.80, indicating modest additional benefit from lightweight personalization. In contrast, oracle style prompts (i.e., nearest annotator and nearest abstract) achieve lower performance, plateauing around 0.75. This may reflect noise introduced during retrieval or label transfer from similar but non-identical contexts, despite their oracle style design. On the test set (Table 1), we use 20 and 50 epochs for training the supervised and unsupervised models, respectively. Our supervised LoRA-tuned models substantially outperform the prior state-of-the-art reported by Guo et al. (2024), where GPT-4 with context learning achieved an F1 score of 0.64 and the oracle classifier reached 0.72. Notably, even our models using vanilla prompting exceed these benchmarks. These results highlight the effectiveness of full supervision combined with parameter-efficient fine-tuning in capturing the semantic relationships between research abstracts and jargon familiarity.
@@
Matadata-enhanced 76.8±1.1 76.1±2.9 77.7±1.0
Profile-enhanced 76.6±1.0 73.7±2.0 79.9±1.6
@@
10% Sup + Unsup 77.0±1.1 78.9±2.6 75.4±0.7
Table 1: Performance of fine-tuned models on the test set. Unless specified, results use vanilla prompting. The $\pm$ values represent standard deviations across 3 repeated runs to illustrate consistency.
Unsupervised fine-tuning without annotations shows limited effectiveness
In the unsupervised setting, models were fine-tuned solely on each annotator’s published papers, without familiarity annotations. Although performance improves slightly over time, overall results remain poor, confirming that publication history alone fails to capture familiarity judgments, consistent with prior findings Haghani (2023). We exclude the nearest annotator and nearest abstract settings from this experiment, as they require access to annotated familiarity labels and are therefore not applicable in the unsupervised scenario.
Unsupervised personalization enables efficient adaptation with minimal supervision.
Our findings show that integrating unsupervised user publication data substantially boosts performance when combined with just 10% of labeled training data, achieving an F1 score of 77.0. This mixed strategy not only approaches the performance of fully supervised models but also significantly outperforms models trained on the same limited labeled data alone. These results underscore the potential of leveraging unlabeled, domain-relevant data to reduce annotation costs while preserving personalization quality. This approach improves the scalability and accessibility of personalized NLP systems, particularly in settings where manual annotation is expensive or infeasible. Additional qualitative analyses, such as model generalization to related tasks, performance across individual annotators, and domain-specific behavior, are in App.§B.
@@ 4 Conclusions
In this work, we addressed the challenge of domain-specific jargon as a barrier to interdisciplinary communication by introducing a novel framework for personalized jargon identification. Leveraging lightweight large language models fine-tuned with LoRA, our approach achieves significant performance gains over prior work while maintaining computational efficiency. We demonstrated the effectiveness of incorporating personalized prompts grounded in a researcher’s background, and showed that integrating unsupervised user data with minimal labeled supervision can nearly match fully supervised performance. Together, these contributions offer a scalable and resource-efficient path toward building personalized NLP tools that enhance accessibility and foster cross-disciplinary understanding.
@@ B.2 Model Improvement Analysis in Terms of Individual Annotators
Taking annotator #4 as the object, a qualitative analysis of the missed and falsely detected jargon reveals several interesting patterns. The baseline model (trained with 1% training set), while showing some capability in jargon detection, struggled with terms that exhibit a combination of characteristics. Firstly, it frequently failed to identify the terms as jargon that are relatively short and composed of common words but carry highly specific meanings within a particular domain. Examples include ‘Radial curves’ (Materials Science), ‘Op-amp’ (Physics), and ‘Domains’ (Geology). These terms, due to their brevity and seemingly ordinary components, may have been harder for the baseline to differentiate from general language use. Secondly, the baseline model had difficulty with multi-word terms where the meaning is not a straightforward combination of the individual words, but rather a more nuanced concept. This is evident in its failure to identify ‘Bayesian optimal mechanism’ (Economics), ‘Riemannian framework’ (Materials Science), ‘Bose-Einstein condensate’ (Physics), ‘Psychometric properties’ (Economics), ‘Dialectical quality’ (Philosophy), ‘Explanatory account’ (Linguistics), ‘Long-range ordered coupling’ (Physics, Materials Science), and ‘Qualitative spatio-temporal inferences’ (Psychology). In these cases, the model may have lacked the ability to capture the semantic relationships and contextual dependencies necessary for accurate identification. Thirdly, the baseline also missed acronyms like ‘CW-SSIM’ (Agricultural And Food Sciences), ‘MANOVA’ (Education), and ‘ARMAX model’ (Business, Engineering). Acronyms often present a challenge due to their condensed nature and lack of explicit semantic clues. Finally, there were instances where the jargon term spans multiple disciplines, such as ‘Monolayers’ (Engineering, Biology), ‘Peri-implant bone density’ (Materials Science, Medicine, Biology), and ‘Regulatory mechanisms’ (Biology, Environmental Science), which might have added to the difficulty. While the improved model demonstrated a higher F1 score, indicative of better overall performance, it exhibited a tendency to produce more false positives. These false positives included terms like ‘Savitzky-Golay (SG) filter’ (Environmental Science), ‘Meta-analyses’ (Medicine), ‘Post-test’ (Education), ‘Quantitative research’ (Education), and ‘Content analysis’ (Medicine). This suggests that the improved model, in its attempt to capture a broader range of jargon, may be more sensitive to terms that share some characteristics with jargon but are more commonly used or understood. This could indicate a trade-off where the improved model sacrifices some precision for increased recall, potentially overgeneralizing in certain contexts. Specifically, the improved model appears to be more prone to misclassifying statistical or methodological terms (e.g., ‘Meta-analyses’, ’Post-test’, ‘Quantitative research’) as jargon, possibly due to their frequent occurrence in academic contexts, even when they are relatively well understood within the broader research community.

### v3

@@ Abstract
Personalizing jargon detection and explanation is essential for making technical documents accessible to readers with diverse disciplinary backgrounds. However, tailoring models to individual users typically requires substantial annotation efforts and computational resources due to user-specific finetuning. To address this, we present a systematic study of personalized jargon detection, focusing on methods that are both efficient and scalable for real-world deployment. We explore two personalization strategies: (1) lightweight finetuning using Low-Rank Adaptation (LoRA) on open-source models, and (2) personalized prompting, which tailors model behavior at inference time without retaining. To reflect realistic constraints, we also investigate semi-supervised approaches that combine limited annotated data with self-supervised learning from users’ publications. Our personalized LoRA model outperforms GPT-4 with contextual prompting by 21.4% in F1 score and exceeds the best performing oracle baseline by 8.3%. Remarkably, our method achieves comparable performance using only 10% of the annotated training data, demonstrating its practicality for resource-constrained settings. Our study offers the first work to systematically explore efficient, low-resource personalization of jargon detection using open-source language models, offering a practical path toward scalable, user-adaptive NLP system  [^1] Code can be found at: anonymous GitHub repository..
@@ 1 Introduction
Figure 1: To enable personalized jargon detection, we first fine tune a LoRA-based PEFT model using supervised, self-supervised, and semi-supervised training strategies that reflect real world scenarios with varying levels of annotation availability. Next, we enhance the contextual understanding of the target researcher through a range of background aware prompting methods, including vanilla, metadata based, profile enhanced, nearest researcher, and nearest abstract), to generate personalized familiarity predictions.
@@
Guo et al. (2024) made the first step toward personalized jargon detection by releasing a benchmark and analyzing GPT-4’s capabilities. However, their approach depends on costly prompting and rich supervision, raising concerns about scalability and generalization. In this paper, we provide the first comprehensive and systematic study of personalized jargon detection with an emphasis on efficiency, scalability, and low-resource practicality. We show that lightweight PEFT on open-source models can surpass GPT-4 while requiring only 10% annotated data, highlighting the feasibility of scalable, user-adaptive NLP systems.
@@ 2 Method
To personalize jargon familiarity, we investigate three fine-tuning settings: supervised, self-supervised, and semi-supervised ( §2.2), and incorporate contextual prompts ( §2.3). To isolate the impact of personalization, we additionally evaluate a leave-one-annotator-out setting ( §2.4).
@@ 2.1 Experimental Setup
We use the personalized jargon detection dataset from Guo et al. (2024), which contains 11k term familiarity (i.e., familiar or unfamiliar) annotations provided by 11 computer science researchers for terms extracted from 100 paper abstracts. To our knowledge, it is the only interdisciplinary jargon dataset with high-quality annotations, rich personal metadata, and accompanying published papers from the annotators. We follow the same data split as Guo et al. (2024), dividing the dataset into 60/20/20 for the train, validation, and test sets. We selected Llama-3.1 8B Instruct 4bit Dubey et al. (2024) as our baseline model for personalized jargon detection based on its strong instruction-following ability, low mismatch rate, and competitive performance in preliminary evaluations across several state-of-the-art LLMs. We evaluate model performance using the effective F-1 Score, which adjusts for output format errors by penalizing predictions with a high mismatch rate. Details of hyperparameters are in App. §A.
2.2 Lightweight Personalized PEFT
  • Supervised: We adopt LoRA (Hu et al., 2022) for PEFT, following findings from Tan et al. (2024b) demonstrating its strong performance. The model is trained to predict annotator familiarity from each term and its associated abstract.
  • Self-supervised: To simulate low-resource scenarios, we fine-tune on unlabeled titles and abstracts from each annotator’s prior publications using a causal language modeling (next-token prediction) objective. For with $\leq 5$ publications, we augment the corpus with papers from their self-defined subdomain, reflecting practical cases where personalized models rely on domain-relevant but unlabeled content.
  • Semi-supervised: We examine a hybrid approach that combines limited labeled data with the annotator’s publication corpus, evaluating how much supervision is needed to balance annotation effort with personalization quality and to guide future annotation strategies.
@@ 2.3 Background-Aware Prompting
Building on prior work Guo et al. (2024); Tan et al. (2024b), we design prompting strategies that add varying levels of researcher-specific context to the model input (full prompts in Table 2), differing in both type and detail of background information:
  • Metadata: Structured features including the annotator’s self-defined subfield (e.g., NLP, computer vision), publication count, average references, year of first publication, and the domain of the current abstract. They serve as lightweight indicators of expertise and familiarity.
  • Profile: Following prior work on user modeling in personalized recommendation Tan et al. (2024b), we use the baseline model to generate a natural language summary of the annotator’s research background based on their metadata.
@@
2.4 Ablation Study
To isolate the effect of personalization, we consider two ablation settings. First, we include a vanilla prompting setup with no additional contextual information (e.g., no metadata, profiles, or nearest-neighbor retrieval). While still personalized, since the model is fine-tuned on annotator specific data (supervised, self-supervised, or semi-supervised), this setting removes auxiliary background features. Second, we evaluate a non-personalized baseline using a leave-one-annotator-out scheme, where the model is trained on data from all annotators except the held-out one. For comparability with the supervised personalized model, we subsample the training data to match the same number of examples, keeping all other parameters identical.
@@ 3 Results
We compare against the best results from Guo et al. (2024), where the oracle uses familiarity ratings from the most similar annotator and GPT-4 prompts include five prior publications. Figure 2(a) shows validation results: supervised models plateau after 20 epochs (reported at 20), while self-supervised models improve more gradually (reported at 50).
Figure 2: Validation performance of supervised (a), self-supervised (b), and semi-supervised (c) models for jargon familiarity detection. Each point is a personalized model fine-tuned for one of 11 annotators with different prompting strategies. GPT-4 and Oracle baselines are best results from Guo et al. (2024), where the Oracle uses ratings from the most similar annotator and GPT-4 prompts include five publications. For semi-supervised models, training steps on the x-axis allow consistent comparison across dataset sizes and reflect relative computational cost.
Supervised fine-tuning outperforms GPT-4 and oracle settings
On the test set (Table 1), vanilla prompting with a supervised personalized model yields the highest F1 (77.9), outperforming GPT-4 with contextual prompting by 21.4% and the oracle baseline by 8.3%. Additional prompting strategies (metadata, profile, nearest annotator, nearest abstract) do not improve over vanilla prompting, suggesting that LoRA fine-tuning itself captures most of the relevant personalized information. The limited gains from these strategies may reflect either noise in background features or insufficient dataset size to reveal their benefits. Overall, these findings highlight the effectiveness of full supervision with PEFT for modeling the link between annotator background knowledge and jargon familiarity, while showing limited added value from more elaborate prompting.
@@
Metadata 76.8±1.1 76.1±2.9 77.7±1.0
Profile 76.6±1.0 73.7±2.0 79.9±1.6
@@
Self + 10% Sup 77.0±1.1 78.9±2.6 75.4±0.7
Leave-one-annotator-out (no personalization)
Sup baseline 64.7±0.8 63.9±1.1 65.5±0.6
   + Metadata $64.3_{\pm 0.9}$ $2.7_{\pm 1.5} 66$$6.1_{\pm 0.3}$
   + Profile $64.9_{\pm 0.8}$ $64.8_{\pm 0.8}$ $65.0_{\pm 0.7}$
Self + 10% Sup 53.4±0.1 47.8±0.2 60.6±0.1
   + Metadata $8.6_{\pm 0.2} 55$$5.2_{\pm 0.3}$ $62.5_{\pm 0.2}$
   + Profile $60.4_{\pm 0.4}$ $64.2_{\pm 0.3}$ $57.0_{\pm 0.4}$
Table 1: Performance of fine-tuned models on the test set. Unless specified, results use vanilla prompting. Oracle setting uses the familiarity ratings from the annotator with the highest agreement on the training set for the annotator. For GPT-4, prompts include five annotator’s publications. The $\pm$ values represent standard deviations across 3 repeated runs to illustrate consistency.
Self-supervised fine-tuning without annotations shows limited effectiveness
In the self-supervised setting, models were fine-tuned solely on each annotator’s published papers, without familiarity annotations. Although performance improves slightly over time, overall results remain poor, confirming that publication history alone fails to capture familiarity judgments, consistent with prior findings Haghani (2023). We exclude the nearest annotator and nearest abstract settings from this experiment, as they require access to annotated familiarity labels and are therefore not applicable in the self-supervised scenario.
Semi-supervised personalization enables efficient adaptation with minimal supervision.
Integrating self-supervised user publication data with only 10% of labeled training data yields an F1 of 77.0, nearly matching fully supervised performance (71.9) and clearly surpassing models trained on the same limited labeled data alone (63.6). This demonstrates the value of leveraging unlabeled, domain-relevant data to reduce annotation costs while preserving personalization quality, improving the scalability and accessibility of personalized NLP systems in settings where manual annotation is costly or infeasible. Additional qualitative analyses, including generalization to related tasks, annotator-specific performance, and domain-specific behavior, are provided in App. §B.
Personalization is necessary to solve the jargon detection task
The personalized model achieves an F1 of 77.9, which is a 20.4% improvement over supervised leave-one-annotator-out testing (64.7) and a 45.9% improvement over the self-supervised + 10% supervised setting (53.4). While background-aware prompting with metadata or profile does not improve performance over vanilla in the personalized setting, it yields gains without personalized annotation: profile improves F1 by 13.1% (60.4 vs. 53.4) and metadata by 9.7% (58.6 vs. 53.4). These results further underscore the importance of personalization for jargon detection and highlight the effectiveness of PEFT under full supervision, with background-aware prompting offering value only when annotations are limited.
@@ 4 Conclusions
In this work, we present a practical and cost-effective approach to personalization in jargon detection. By fine-tuning lightweight language models with LoRA, our method achieves significant performance gains over prior work while maintaining computational efficiency. We further show that personalized prompts grounded in a researcher’s background improve non-personalized models for familiarity prediction, providing an alternative when direct annotations are unavailable. Remarkably, our method achieves comparable performance with only 10% of annotated data, underscoring its practicality in resource-constrained settings where large-scale annotation is costly or infeasible. Together, these contributions demonstrate the effectiveness and scalability of personalized NLP, offering a path toward tools that improve accessibility and foster cross-disciplinary collaboration.
@@ Appendix A Setup
Implementation Details
We fine-tuned the unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit model with a maximum sequence length of 2048 tokens. For parameter-efficient training, we applied LoRA with rank 16, scaling factor ($\alpha$) 16, dropout 0, and targeted the projection modules (q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj). The model was trained for 100 epochs with a per-device batch size of 2 and gradient accumulation of 4 steps, using a learning rate of 2e-4, weight decay of 0.01, and the AdamW (8-bit) optimizer with a linear scheduler and 5 warmup steps. Training was conducted in FP16 or BF16 precision depending on hardware support. Checkpoints were saved every epoch_size * 10 // 8 steps (approximately every 10 epochs).
@@ B.2 Model Improvement Analysis in Terms of Individual Annotators
Taking annotator #4 as the object, a qualitative analysis of the missed and falsely detected jargon reveals several interesting patterns. The baseline model (trained with 1% training set), while showing some capability in jargon detection, struggled with terms that exhibit a combination of characteristics. Firstly, it frequently failed to identify the terms as jargon that are relatively short and composed of common words but carry highly specific meanings within a particular domain. Examples include ‘Radial curves’ (Materials Science), ‘Op-amp’ (Physics), and ‘Domains’ (Geology). These terms, due to their brevity and seemingly ordinary components, may have been harder for the baseline to differentiate from general language use. Secondly, the baseline model had difficulty with multi-word terms where the meaning is not a straightforward combination of the individual words, but rather a more nuanced concept. This is evident in its failure to identify ‘Bayesian optimal mechanism’ (Economics), ‘Riemannian framework’ (Materials Science), ‘Bose-Einstein condensate’ (Physics), ‘Psychometric properties’ (Economics), ‘Dialectical quality’ (Philosophy), ‘Explanatory account’ (Linguistics), ‘Long-range ordered coupling’ (Physics, Materials Science), and ‘Qualitative spatio-temporal inferences’ (Psychology). In these cases, the model may have lacked the ability to capture the semantic relationships and contextual dependencies necessary for accurate identification. Thirdly, the baseline also missed acronyms like ‘CW-SSIM’ (Agricultural And Food Sciences), ‘MANOVA’ (Education), and ‘ARMAX model’ (Business, Engineering). Acronyms often present a challenge due to their condensed nature and lack of explicit semantic clues. Finally, there were instances where the jargon term spans multiple disciplines, such as ‘Monolayers’ (Engineering, Biology), ‘Peri-implant bone density’ (Materials Science, Medicine, Biology), and ‘Regulatory mechanisms’ (Biology, Environmental Science), which might have added to the difficulty. While the improved model demonstrated a higher F1 score, indicative of better overall performance, it exhibited a tendency to produce more false positives. These false positives included terms like ‘Savitzky-Golay (SG) filter’ (Environmental Science), ‘Meta-analyses’ (Medicine), ‘Post-test’ (Education), ‘Quantitative research’ (Education), and ‘Content analysis’ (Medicine). This suggests that the improved model, in its attempt to capture a broader range of jargon, may be more sensitive to terms that share some characteristics with jargon but are more commonly used or understood. This could indicate a trade-off where the improved model sacrifices some precision for increased recall, potentially overgeneralizing in certain contexts. Specifically, the improved model appears to be more prone to misclassifying statistical or methodological terms (e.g., ‘Meta-analyses’, ‘Post-test’, ‘Quantitative research’) as jargon, possibly due to their frequent occurrence in academic contexts, even when they are relatively well understood within the broader research community.